OpenStudy (anonymous):

Evaluate the function for y the given value of x. f(x)=5x-, if x<-2 x-9, if x less than or greater to -2 f(-2)

OpenStudy (anonymous):

Are you sure that's how the problem is given, because practically what you're saying is:\[f(x) = 5x-\]if\[x<-2\]and \[f(x)=x-9\]if\[x <> -2\]Did you maybe mean on the second choice x greater than or equal to -2?

OpenStudy (anonymous):

yes sometimes i get confused

OpenStudy (anonymous):

lol.. no worries. Well if that's the case then it means that if x=-2, you need to evaluate the function somewhere where it is defined like that. Since the function is saying f(x)=x-9 IF x is greater or equal to -2, then you need to evaluate the function at that expression. Try that, and show me the result you got.

OpenStudy (anonymous):

-11 for both

OpenStudy (anonymous):

Yes, but you don't need to evaluate both. You just need to evaluate f(x)=x-9... But you're correct anyhow.

OpenStudy (anonymous):

ohh okay i see now so its -11?

OpenStudy (anonymous):

Yeah, its -11

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

ur welcome!